OpenStudy (anonymous):

solve the inequality 2/3 (t+2)le 1/4(2t-6)

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } (t+2) \le \frac{ 1 }{ 4 } (2t-6)\]

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

can you simplify? Do you know the distributive property?

OpenStudy (anonymous):

I do but I was under the impression that I need to clear the fractions first, but every time I do that it doesn't match the answer to the back of the book so I'm concerned I'm clearing the fractions incorrectly.